# IntGrad MT: Eliciting LLMs’ Machine Translation Capabilities with Sentence Interpolation and Gradual MT

### v2

@@ Appendix G Full Results of Ablation Study
Table 6 shows the results for every combination of strategies that we explored in the ablation study(§6.1) with En-Ko translation. Table 7 presents the average results and changes in xCOMET scores for each combination of start selection strategies, the number of start sentences, and MT aggregation strategies.
@@
Thr. - 90.65 91.30
Delta - 92.10 92.46
Thr. + Delta - 92.06 92.01
Table 7: Average xCOMET scores and score changes of selected outputs when applying every combination of strategies for start selection and aggregation in En-Ko translation. ‘Score change’ is calculated only for the adopted outputs. ‘All’ selects every output. ‘Thr.’ applies zero-shot translation first and uses IntGrad MT only when the CometKiwi score is below a threshold. ‘Delta’ selects outputs only if they outperform zero-shot translations. ‘Thr. + Delta’ combines these two strategies. We also report the number of end sentences for which interpolation and Gradual MT is executed (‘No. of Interpolated End Sents’) and the number of end sentences for which the IntGrad MT output is selected over zero-shot translation (‘No. of Selected Outputs’). The results show that ‘Thr. + Delta’ reduces the number of interpolation by more than half while maintaining nearly the same translation performance.
Output Adoption Strategy Avg. Score Avg. Score Change of Selected Outputs No. of Interpolated Sents (%) No. of Selected Outputs (%)
@@
Thr. 91.01 3.70 10468 (48%) 10468 (48%)
Delta 92.21 5.30 21864 (100%) 11270 (52%)
Thr. + Delta 91.81 7.98 10468 (48%) 6664 (30%)

### v3

@@ Appendix G Results from Ablation
Table 17 shows the results for every combination of strategies that we explored in the ablation study(6) with En-Ko translation. Table 18 presents the average results and changes in xCOMET scores for each combination of start selection strategies, the number of start sentences, and MT aggregation strategies.
@@
Pre - 90.65 91.30
Post - 92.10 92.46
Pre&Post - 92.06 92.01
Table 18: Average xCOMET scores and score changes of selected outputs when applying every combination of strategies for start selection and aggregation in En-Ko translation. ‘Score change’ is calculated only for the adopted outputs. ‘All’ selects every output. ‘Pre-filtering’, denoted as ‘Pre’, applies zero-shot translation first and uses BridG MT only when the CometKiwi score is below a threshold. ‘Post-filtering’, denoted ans ‘Post’, selects outputs only if they outperform zero-shot translations. ‘Pre- & Post- filtering’, denoted as ‘Pre & Post’ combines these two strategies. We also report the number of end sentences for which bridging and Gradual MT is executed (‘No. of Interpolated End Sents’) and the number of end sentences for which the BridG MT output is selected over zero-shot translation (‘No. of Selected Outputs’). The results show that ‘Pre & Post’ reduces the number of bridging by more than half while maintaining nearly the same translation performance.
Filtering Strategy Avg. Score Avg. Score Change of Selected Outputs No. of Interpolated End Sents(%) No. of Selected Outputs (%)
@@
Pre 91.01 3.70 10468 (48%) 10468 (48%)
Post 92.21 5.30 21864 (100%) 11270 (52%)
Pre&Post 91.81 7.98 10468 (48%) 6664 (30%)
Appendix H Evaluation Results of BridG MT with Varying Start Pool Quality
Figure 5 shows xCOMET scores of BridG MT when using start pools of varying quality.
Figure 5: xCOMET scores of BridG MT when using start pools of varying quality. In each plot, the horizontal axis represents the xCOMET score of the start pool, while the vertical axis represents the xCOMET score of the final output. A single start sentence was used for evaluation, without the application of any filtering method.
Appendix I Analysis On Sentence Bridging
Below we present the analysis we conducted on sentence bridges. We analyzed sentence bridges from En-Ko translation results.
I.1 Progresses of Bridging.
We examined whether LLMs genuinely bridge start and end sentences or simply generate random sentences. To assess this, we use SBERT to embed the sentences in bridge and calculate their Euclidean distances from the end sentence. If these distances generally decrease, it indicates successful bridging. To mesure this, we defined progress as $\text{progress}=d_{n-1,e}-d_{n,e}$ where $d_{n,e}$ denotes the euclidian distance between $n$th sentence in each bridge and end sentence.
Figure 6: Distribution of average progress from each sentence bridge. The sentence bridges are drawn from the English-to-Korean BridG MT results. Progress indicates how much each sentence in the bridge moves closer to the target sentence in terms of Euclidean distance.
As shown in Figure 6, the average progress of each bridge is generally positive on both the source(English) and target(Korean) sides, indicating that the sentence bridging effectively connects the two sentences. The mean and standard deviation of average progress across bridges were 0.12 and 0.27 for English, and 0.06 and 0.031 for Korean, respectively.
I.2 Visualization of Bridges
In addition to the analysis on progress, we conducted visualizations on both the source and target sides to examine whether our experimental results align with the intuition illustrated in Figure 2. Specifically, we visualized the embeddings of the sentences within the sentence bridges and the Korean translations generated at each step of Gradual MT. For English sentences, we used all-mpnet-base-v2 to obtain embeddings, while for Korean sentences, we used intfloat/multilingual-e5 (Wang et al., 2024). On the Korean side, we additionally plotted 50-shot example translations to examine whether the intermediate translations produced at each Gradual MT step gradually move closer to the embedding of the gold translation, and how they compare to the few-shot examples. The results of this analysis show that sentence bridging effectively generates intermediate sentences that bridge the start and end sentences on the source side. Furthermore, we frequently observed that the intermediate translations from Gradual MT also move progressively closer to the gold translation. These findings support the intuition underlying our approach. The visualizations are shown in Figure 7 and Figure 8.
Figure 7: 24 samples of 2D scatter plots of embeddings from sentence bridges. The plots are projected from SBERT embeddings onto a 2D plane using PCA. The X and Y axes of each plot represent the first and second principal components, respectively. Arrows in each plot show the trajectory of sentence shifts from the start sentence (blue-colored dot) to the end sentence (yellow-colored dot).
Figure 8: 12 2D scatter plots of embeddings from Gradual MT. Sentence embeddings from the intfloat/multilingual-e5-large model are projected onto a 2D plane using PCA. The X and Y axes represent the first and second principal components, respectively. Green dots and sky-blue triangles indicate the few-shot examples and 50-shot MT outputs. Purple and blue X marks denote the translations from each step and the final translation generated by Gradual MT.